OpenStudy (anonymous):

Can you help I want to know the steps in proving it. Verify the identity. cotx minus pi divided by two. = -tan x

OpenStudy (anonymous):

\[\cot \left[ x-\left(\begin{matrix}\Pi \\ 2\end{matrix}\right) \right]\]

OpenStudy (anonymous):

= - tan x